# A new Angle on Benchmarking Noncovalent Interactions

### v2

@@ 3.4 Local Correlation
Having established the basis set effects, we can estimate canonical CCSD(T) at the basis set limit. Furthermore, especially with the advent of using local correlation methods and extrapolating them,69, 70, 71, 72 we can evaluate their accuracy when going to larger molecules. For this purpose, we compare these methods, especially for a small basis set of double-zeta quality. Again, all of the investigated series exhibit nearly perfect linear regression with an value of 0.996 for methods.
@@
However, there is one caveat: When increasing the LNO cutoff settings Normal Tight VeryTight vvTight (‘very very tight’), unlike for basis sets, the slopes do not always converge to the canonical value for local CCSD(T). This makes recently promoted CPS extrapolations more erratic than, e.g., extrapolations to the complete basis set (CBS) limit. Basis sets are constructed to exhibit a monotonic convergence with increasing size to the basis set limit. In contrast, LNO and DLPNO cutoffs just define which natural orbitals are neglected- the energies do not have to increase or decrease monotonically. For example, for the cc-pVDZ basis set is applied, the differences in slopes when going from Normal to Tight to VeryTight to vvTight LNO cutoffs to canonical CCSD(T) change by 2.1, 0.8, 1.0, and 0.7 kJ/subunit for the cc-pVDZ basis set and 2.3, 0.8, 0.7, 1.3 kJ/subunit for the aug-cc-pVTZ basis set, respectively. For the latter basis set, we thus obtain a gap from the VeryTight to the vvTight setting and canonical CCSD(T). From these numbers, if we would just consider LNO-CCSD(T), we would be bound to believe that the slopes are close to converging, as we would get a slope change of 2.3, 0.8, and 0.7 kJ/subunit . This is, however, not the case, as the difference to the canonical slope is not less than 0.7 kJ/subunit .
This indicates that the CPS extrapolation, especially like in our context, should be investigated further for larger molecules, and its convergence may not be as strict as warranted.
@@
DLPNO, while being rather close to the VeryTight and vvTight LNO-CCSD(T) cutoff, sometimes even moves into the wrong direction and is increasing its slope further when the PNO-cutoff is tightened. Interestingly, this seems to be rather basis set dependent, with the basis sets without diffuse functions (cc-pVTZ and cc-pVQZ) behaving more like the cc-pVDZ basis set. Here, an extrapolation does not work — while for the aug-cc-pVTZ basis set, the CPS extrapolation is close to the canonical CCSD(T) value.
Although the differences in slopes when going from vvTight LNO-CCSD(T) to canonical CCSD(T) are somewhat larger than e.g. when going from an aug-cc-pVTZ basis set to the CBS limit, these differences are rather small. Even for slopes of larger molecules, the errors are still minor when using LNO-CCSD(T) and/or a basis set of limited size.
For the ethylene stack series the differences in slopes when going from Normal to Tight to VeryTight to vvTight to canonical CCSD(T) change by 0.78, 0.48, -0.01, and 0.19 kJ/subunit, which also implies that the CPS extrapolation would only work for rather loose convergence criteria.
@@
In this series, it is noteworthy that the DLPNO-CCSD(T) method with a TCutPNO value of slightly underestimates the slope but yields results that are closest to those obtained with the canonical CCSD(T) method. In contrast, the DLPNO-CCSD(T) method with TCutPNO set to overestimates the slope relative to the canonical CCSD(T), yet it still provides a more accurate slope estimate than any of the LNO methods. Furthermore, the application of CPS extrapolation produces values that are very close to those obtained with the LNO-CCSD(T) method using a vvTight threshold. However, it does not improve upon the accuracy achieved with the TCutPNO setting of .
@@ 3.5 Electronic Structure Methods
Here, we investigate several post-Hartree-Fock methods, such as MP2, MP2.5, MP3, the dispersion contribution to DFT-SAPT, as well as dRPA. From these methods, DFT-SAPT and dRPA in particular are deemed to be as more accurate than perturbation theory for large molecules.11, 12 For the acene series, we obtain the trend of MP3 CCSD DFT-SAPT dRPA CCSD(T) MP2.5 MP2. This is hardly surprising, as MP2 very much overestimates the interaction energies of large aromatic molecules, while MP3 underestimated them, as does CCSD. Of course, the question arises about the accuracy of CCSD(T), dRPA, and DFT-SAPT.
For the ethylene stack series, we obtain the trend CCSD MP3 DFT-SAPT dRPA CCSD(T) MP2.5 MP2. As observed in the acene series, MP2 overestimates dispersion contributions while both CCSD and MP3 underestimate them. dRPA yields results similar to CCSD(T), which is clearly observed for the cc-pVZ (X=D,T,Q) basis sets. (We note73 that RPA is equivalent to a ring-diagram simplification of coupled cluster with all doubles, CCD.)
@@ 3.6 Effects of the Geometry
So far, we have used geometries for both acenes and ethylene stacks in which the distance between the two monomers were fully optimized within its (high) symmetry. In order to show that the linear behavior is not an artefact, it may be worth to investigate another geometry with a different intermonomer distance. Here, we chose the stacks at the rather large optimized distance of the ethylene dimer, as shorter distances (e.g. at the hexatetraene distance) would result in a positive interaction energy for the ethylene dimer. The ethylene stack series with relaxed geometry exhibits a slope of 8.70 0.27 kJ/subunit. In contrast, the same series with the fixed geometry at the minimum ethylene dimer distance of 4.46 Å shows a much smaller slope of 1.26 0.24 kJ/subunit, where the uncertainty is about one-fifth of the actual value.
@@
All methods analyzed in this study demonstrate a better convergence of the LNO approaches compared to canonical CCSD(T) for relaxed geometries, irrespective of the basis set used. This is attributed to the fact that for these geometries, the systems energy is closer to the true potential energy minimum for the respective method and basis set. This conclusion is supported by the observation that the absolute values of total energies and correlation energies are greater. Similarly, basis set convergence is more consistent for relaxed geometries, as the electron distribution and correlation effects are more accurately aligned with the optimized structure, resulting in more consistent convergence behavior as the basis set is expanded.
@@ 3.7 Best Estimates of Slopes
Usually, the best estimates in energies are obtained by going to ever larger basis sets and then just extrapolating the energies to the complete basis set (CBS) limit. In our case, however, we can either calculate ever larger systems along the sequences studied with a smaller basis set or increase the basis set size thus adding one dimension to the problem. Our results are summarized in Tables 2 and 3. The slopes obtained from just the first two acene dimers, namely benzene2 and naphthalene2, barely differ compared to those of the whole series up to pentacene; the cp-corrected DZ CCSD(T) slope changes by 0.26 kJ/subunit, and the cp-uncorrected DZ CCSD(T) slope even less at 0.17 kJ/subunit. In general, including larger systems in the fits increases the slopes. From these values, including the cc-pV{T,Q}Z basis set extrapolation which is still feasible for canonical CCSD(T) for the benzene and naphthalene dimer, we can estimate the canonical CCSD(T)/cbs values for the benzene-pentacene series, which will be -23.14-0.26=-23.40 kJ/subunit for the cp-corrected CCSD(T) results and -23.09-0.17=-23.26 kJ/subunit for cp-uncorrected CCSD(T). Finally, this values can then be extended to the hexacene dimer with an aug-cc-pV{Q,5}Z basis set extrapolation.
@@
Of more interest are the post-CCSD(T) slopes. The CCSDT and CCSDT(Q) results for acene dimers (benzene and naphthalene dimers) included in Tables 2 and 3 were obtained using rank-reduced coupled cluster methods described in Refs. 74, 75, 76. In these calculations, the coupled cluster equations are solved within a certain excitation subspace instead of the full space as in the canonical methods. This allows for reducing the costs of the calculations, but introduces an additional parameter (size of the excitation subspace) as a parameter in the rank-reduced formalism. Therefore, it is important that the results are stable with respect to the adopted value of this parameter. In the Supporting Information, we provide a detailed analysis of the convergence of the rank-reduced results with respect to size of the excitation subspace for benzene dimer and naphthalene dimer, showing that the numerical values reported here are sufficiently well converged for the present purposes. The only additional approximation employed in the rank-reduced calculations is the Cholesky decomposition of two-electron integrals. We used the full pivoting variant of this decomposition with a VeryTight threshold () for the diagonal elements. The errors resulting from the decomposition with these settings impact the calculated interaction energies by no more than a few thousands of kJ/mol. The CCSDT method reduces the slopes (in absolute value) by a sizeable 0.75 (cp-corrected DZ) and 0.78 (standard DZ) kJ/subunit. However, CCSDT(Q) (again, in absolute value) reverts this by 0.4 in both instances, shrinking the CCSD(T) to CCSDT(Q) differences in the slopes to 0.35 (cp-corrected DZ) and 0.38 (uncorrected DZ) kJ/subunit.
@@
(We note in passing that the (Q) contribution found here for benzene dimer, 0.50 kJ/mol, is just over half of the value of Karton and Martin77, 0.187–0.200 kcal/mol, or 0.78–0.84 kJ/mol. The latter was obtained from a thermochemical cycle combined with a CCSDT(Q) calculation in an unpolarized double-zeta basis set, and serves as a cautionary tale in this regard.)
@@
For the Hartree-Fock exchange part, this is rather simple, as all calculations were performed at the largest basis sets and the cp-corrected and cp-uncorrected slopes are 11.95 and 11.96 kJ/subunit, respectively, yielding exchange-correlation slopes of -10.5 0.2 kJ/mol per acene ring dimer. For the largest acene dimer hexacene, this would make an interaction energy of approximately 63.0 kJ/mol, considering our best estimates. This value, however, is neglecting the intercepts of 3.3 for Hartree-Fock exchange and 1.9 for correlation, arriving at 57.8 1.4 kJ/mol. CCSD(T), for comparison, would yield slightly larger slopes of -11.1 and -10.9, with an interaction energy of 66.0 for the hexacene dimer. Including the intercepts (3.3 and 2.0), we arrive at a final CCSD(T) energy of 60.7 1.5 kJ/mol.
| Method | |||||||||
| CCSD(T) | post-CCSD(T) | ||||||||
| Slopes | Basis Set | MP2 | CCSD | VeryTight | vvTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -6.53 | -4.74 | -5.56 | -5.58 | -5.47 | -5.35 | -5.36 | -5.49 |
| 13 | DZ | -6.99 | -4.89 | -5.83 | -5.81 | -5.69 | -5.56 | -5.54 | -5.67 |
| 14 | DZ | -7.44 | -5.08 | -6.12 | -6.07 | -5.95 | -5.81 | -5.79 | -5.92 |
| 12 | TZ | -8.40 | -5.82 | -7.01 | -7.05 | -6.95 | -6.81∗ | -6.80∗ | -6.93∗ |
| 17 | TZ | -9.93 | -6.38 | -7.95 | -7.96 | -7.77 | -7.62∗ | -7.61∗ | -7.74∗ |
| 12 | QZ | -9.06 | -6.24 | -7.63 | -7.67 | -7.54 | -7.40∗ | -7.39∗ | -7.52∗ |
| 15 | QZ | -10.33 | -6.72 | -8.40 | -8.42 | -8.24 | -8.10∗ | -8.08∗ | -8.21∗ |
| 17 | QZ | -8.54 | -8.58 | -8.37 | -8.23∗ | -8.21∗ | -8.35∗ | ||
| 12 | a{T,Q}Z | -9.53 | -6.55 | -8.04 | -8.07 | -7.98 | -7.84∗ | -7.82∗ | -7.96∗ |
| 17 | a{T,Q}Z | -8.79 | -8.65∗ | -8.64∗ | -8.77∗ | ||||
| 12 | a{Q,5}Z | -9.54 | -6.55 | -8.07 | -8.04 | -7.98 | -7.84∗ | -7.83∗ | -7.96∗ |
| 15 | a{Q,5}Z | -8.67 | -8.53∗ | -8.52∗ | -8.65∗ | ||||
| 17 | a{Q,5}Z | -8.79 | -8.65∗ | -8.64∗ | -8.77∗ | ||||
@@
| Method | |||||||||
| CCSD(T) | post-CCSD(T) | ||||||||
| Slopes | Basis Set | MP2 | CCSD | VeryTight | vvTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -7.03 | -4.87 | -5.91 | -5.93 | -5.72 | -5.59 | -5.59 | -5.72 |
| 13 | DZ | -7.60 | -5.15 | -6.26 | -6.24 | -6.09 | -5.94 | -5.92 | -6.06 |
| 14 | DZ | -8.14 | -5.44 | -6.66 | -6.62 | -6.46 | -6.30 | -6.28 | -6.41 |
| 12 | TZ | -9.09 | -6.27 | -7.61 | -7.64 | -7.53 | -7.38∗ | -7.36∗ | -7.49∗ |
| 17 | TZ | -10.71 | -6.94 | -8.62 | -8.62 | -8.46 | -8.30∗ | -8.28∗ | -8.42∗ |
| 12 | QZ | -9.41 | -6.45 | -7.91 | -7.93 | -7.81 | -7.65∗ | -7.63∗ | -7.76∗ |
| 15 | QZ | -10.72 | -6.97 | -8.71 | -8.72 | -8.55 | -8.39∗ | -8.37∗ | -8.51∗ |
| 17 | QZ | -8.83 | -8.85 | -8.67 | -8.51∗ | -8.49∗ | -8.63∗ | ||
| 12 | a{T,Q}Z | -9.44 | -6.30 | -7.78 | -7.79 | -7.72 | -7.56∗ | -7.54∗ | -7.68∗ |
| 17 | a{T,Q}Z | -8.65 | -8.49∗ | -8.47∗ | -8.60∗ | ||||
| 12 | a{Q,5}Z | -9.58 | -6.53 | -8.04 | -8.04 | -7.95 | -7.80∗ | -7.78∗ | -7.91∗ |
| 15 | a{Q,5}Z | -8.69 | -8.54∗ | -8.52∗ | -8.65∗ | ||||
| 17 | a{Q,5}Z | -8.88 | -8.72∗ | -8.70∗ | -8.84∗ | ||||
@@
An analogous estimation is performed for the ethylene stack series in Table 4 and 5. To extrapolate CCSDT and CCSDT(Q) values, we utilize the difference between CCSDT and CCSD(T) or between CCSDT(Q) and CCSD(T) results obtained with the cc-pVDZ basis set. The most accurate CCSD(T) calculations were performed using the cc-pVTZ basis set for the entire ethylene stack series (ethylene, tetraene, pentaene up to tetradecaneheptaene) and the cc-pVQZ basis set for ethylene - decanepentaene. These results were subsequently used to extrapolate values to larger basis sets. Interestingly, canonical CCSD(T) is even slightly underestimating the correlation interaction energy slopes relative to CCSDT(Q), showing that the correlation slopes obtained are rather dependent on the systems chosen. Overall, the post-CCSD(T) contributions are also smaller for the ethylene stacks, as the CCSDT-CCSD(T) difference amounts to 2.7% of the correlation contribution to the interaction energy. For the acene series, the CCSDT-CCSD(T) difference is more substantial at 4.6% — presumably because of the aromatic rings.
The more economical CCSD(T)Λ method yields qualitatively the same answer as CCSDT-3 and CCSDT-2, but not quantitatively.

### v3

@@ 3.4 Local Correlation
Having established the effects of the basis set, we can estimate canonical CCSD(T) at the basis set limit. Furthermore, especially with the advent of local correlation methods and extrapolating in terms of their cutoff parameters,71, 72, 73, 74 we can evaluate their accuracy when going to larger molecules. For this purpose, we compare these methods, especially for a small basis set of double-zeta quality. Again, all of the investigated series exhibit nearly perfect linear regression with values of 0.996 for all methods.
@@
However, there is one caveat: When increasing the LNO cutoff settings Normal Tight VeryTight vvTight (‘very very tight’), unlike with basis sets, the slopes do not always converge to the canonical value for local CCSD(T). This makes recently promoted complete PNO space (CPS 75) extrapolations more erratic than, e.g., extrapolations to the complete basis set (CBS) limit. Correlation consistent53 basis sets are constructed to exhibit smooth and monotonic convergence to the complete basis set limit. In contrast, LNO and DLPNO cutoffs simply define which natural orbitals are neglected- the energies do not have to increase or decrease monotonically. For example, when the DZ basis set is applied, the differences in slopes when going from Normal to Tight to VeryTight to vvTight LNO cutoffs to canonical CCSD(T) change by 2.1, 0.8, 1.0, and 0.7 kJ per subunit for the DZ basis set and 2.3, 0.8, 0.7, 1.3 kJ per subunit for the aTZ basis set, respectively. For the latter basis set, we thus obtain a gap from the VeryTight to the vvTight setting and canonical CCSD(T). From these numbers, if we would just consider LNO-CCSD(T), we would be bound to believe that the slopes are close to converging, as we would get a slope change of 2.3, 0.8, and 0.7 kJ per subunit. This is, however, not the case, as the difference to the canonical slope remains no less than 0.7 kJ per subunit.
@@
This indicates that the CPS extrapolation, especially like in our context, should be investigated further for larger molecules, and its convergence may not be as strict as desirable. DLPNO, while generally rather close to LNO-CCSD(T) with VeryTight and vvTight cutoffs, sometimes even moves into the wrong direction — increasing its slope further as the PNO-cutoff is tightened. Interestingly, this seems to be rather basis set-dependent, with the basis sets without diffuse functions (TZ and QZ) behaving more like the DZ basis set. For these, an extrapolation does not work — while for the aTZ basis set, the CPS extrapolation is close to the canonical CCSD(T) value. Although the differences in slopes when going from vvTight LNO-CCSD(T) to canonical CCSD(T) are somewhat larger than e.g. from an aTZ basis set to the CBS limit, these differences are relatively small. Even for slopes of larger molecules, the errors are still mild when using LNO-CCSD(T) and/or a basis set of limited size.
@@
For the ethylene stack series, the differences in slopes when going from Normal to Tight to VeryTight to vvTight to canonical CCSD(T) change by 0.78, 0.48, -0.01, and 0.19 kJ per subunit, which also implies that the CPS extrapolation would only work for rather loose convergence criteria. In this series, it is noteworthy that the DLPNO-CCSD(T) method with a TCutPNO value of slightly underestimates the slope but yields results that are closest to those obtained with the canonical CCSD(T) method. In contrast, the DLPNO-CCSD(T) method with TCutPNO set to overestimates the slope relative to the canonical CCSD(T), yet it still provides a more accurate slope estimate than any of the LNO methods. Furthermore, the application of CPS extrapolation produces values that are very close to those obtained with the LNO-CCSD(T) method using a vvTight threshold. However, it does not improve upon the accuracy achieved with the TCutPNO setting of .
@@ 3.5 Electronic Structure Methods
Here, we investigate several post-Hartree-Fock methods, such as MP2, MP2.5, MP3, the dispersion contribution to DFT-SAPT, as well as dRPA. Among these methods, DFT-SAPT and dRPA in particular are deemed to be more accurate than perturbation theory for large molecules.11, 12 For the acene series, we obtain the trend of MP3 CCSD DFT-SAPT dRPA CCSD(T) MP2.5 MP2. This is hardly surprising, as MP2 significantly overestimates the interaction energies of large aromatic molecules, while MP3 underestimated them, as does CCSD. Of course, the question arises about the accuracy of CCSD(T), dRPA, and DFT-SAPT.
For the ethylene stack series, we observe the trend CCSD MP3 DFT-SAPT dRPA CCSD(T) MP2.5 MP2. As with the acene series, MP2 overestimates dispersion contributions while both CCSD and MP3 underestimate them. dRPA yields results similar to CCSD(T), as observed with the cc-pVZ (X=D,T,Q) basis sets. (We note76 that RPA is equivalent to a ring-diagram simplification of coupled cluster with all doubles, CCD.)
@@ 3.6 Effects of the Geometry
So far, we have used geometries for both acenes and ethylene stacks in which the distances between the two monomers were fully optimized within the (high) symmetry imposed. To show that the linear behavior is not an artefact, it is worth investigating another geometry with a different intermonomer distance. Here, we chose the stacks at the rather large optimized intermonomer distance of the ethylene dimer, as shorter distances (e.g., at the hexatetraene distance) would result in a positive interaction energy for the ethylene dimer. The ethylene stack series with relaxed geometry exhibits a slope of -8.72 0.27 kJ per subunit. In contrast, the same series with fixed geometry at the minimum ethylene dimer distance of 4.46 Å shows a much smaller slope of -3.81 0.24 kJ per subunit.
@@
All methods analyzed in this study demonstrate better convergence of the LNO approaches compared to canonical CCSD(T) for relaxed geometries, irrespective of the basis set used. This is attributed to the fact that, for these geometries, the system’s energy is closer to the true potential energy minimum. This conclusion is supported by the observation that the absolute values of total energies and correlation energies are larger. Similarly, basis set convergence is more consistent for relaxed geometries, as the electron distribution and correlation effects are more accurately aligned with the optimized structure, resulting in more consistent convergence behavior as the basis set is expanded.
@@ 3.7 Best Estimates of Slopes
Usually, the best estimates in energies are obtained by going to ever larger basis sets and then just extrapolating the energies to the complete basis set (CBS) limit. In our case, however, we can either calculate ever larger systems along the sequences studied with a smaller basis set or increase the basis set size thus adding one dimension to the problem. Our results are summarized in Tables 2 and 3. The slopes obtained from just the first two acene dimers, namely benzene2 and naphthalene2, barely differ compared to those of the whole series up to pentacene; the CP-corrected DZ CCSD(T) slope changes by 0.26 kJ per subunit, and the CP-uncorrected DZ CCSD(T) slope even less at 0.17 kJ per subunit. In general, including larger systems in the fits increases the slopes. From these values, including the cc-pV{T,Q}Z basis set extrapolation, which is still feasible for canonical CCSD(T) for the benzene and naphthalene dimers, we can estimate the canonical CCSD(T)/cbs values for the benzene-pentacene series, which will be kJ per subunit for the CP-corrected CCSD(T) results and kJ per subunit for CP-uncorrected CCSD(T). Finally, these values can then be extended to the hexacene dimer with an aug-cc-pV{Q,5}Z basis set extrapolation by using the LNO-CCSD(T) values, e.g. arriving at -23.40-24.55-(-24.89)=-23.07 for the CP-corrected CCSD(T) estimate.
@@
Of more interest are the post-CCSD(T) slopes. The CCSDT and CCSDT(Q) results for acene dimers (benzene2 and naphthalene2) included in Tables 2 and 3 were obtained using rank-reduced coupled cluster methods described in Refs. 77, 78, 79. In these calculations, the coupled cluster equations are solved within a certain excitation subspace instead of the full space as in the canonical methods. This allows for reducing the costs of the calculations, but introduces an additional parameter (size of the excitation subspace) as a parameter in the rank-reduced formalism. Therefore, it is important that the results are stable with respect to the adopted value of this parameter. In the Supporting Information, we provide a detailed analysis of the convergence of the rank-reduced results with respect to the size of the excitation subspace for benzene dimer and naphthalene dimers, showing that the numerical values reported here are sufficiently well converged for the present purposes. The only additional approximation employed in the rank-reduced calculations is the Cholesky decomposition of two-electron integrals. We used the full pivoting variant of this decomposition with a VeryTight threshold () for the diagonal elements. The errors resulting from the decomposition with these settings impact the calculated interaction energies by no more than a few thousandths of kJ/mol. The CCSDT method reduces the slopes (in absolute value) by a sizeable 0.75 (CP-corrected DZ) and 0.78 (standard DZ) kJ per subunit. However, CCSDT(Q) (again, in absolute value) reverts this by 0.4 in both instances, shrinking the CCSD(T) to CCSDT(Q) differences in the slopes to 0.35 (CP-corrected DZ) and 0.38 (uncorrected DZ) kJ per subunit.
@@
(We note in passing that the (Q) contribution found here for benzene dimer, 0.50 kJ/mol, is just over half of the value reported by Karton and Martin80, 0.19–0.20 kcal/mol, or 0.78–0.84 kJ/mol. This was obtained from a thermochemical cycle combined with a CCSDT(Q) calculation in an unpolarized cc-pVDZ(p,s) double-zeta basis set, i.e. [3s2p] on C and [2s] on H, and serves as a cautionary tale in this regard — clearly, one neglects the functions on carbon at one’s peril.)
@@
For the Hartree-Fock exchange part, this is rather simple, as all calculations were performed at the largest basis sets and the CP-corrected and CP-uncorrected slopes are 11.95 and 11.96 kJ per subunit, respectively, yielding exchange-correlation slopes of -10.5 0.2 kJ/mol per acene ring dimer. For the largest acene dimer hexacene, this would make an interaction energy of approximately 63.0 kJ/mol, considering our best estimates. This value, however, is neglecting the intercepts of 3.3 for Hartree-Fock exchange and 1.9 for correlation, arriving at 57.8 1.4 kJ/mol. CCSD(T), for comparison, would yield slightly larger slopes of -11.1 and -10.9 kJ/mol, with an interaction energy of 66.0 kJ/mol for the hexacene dimer. Including the intercepts (3.3 and 2.0 kJ/mol), we arrive at a final CCSD(T) energy of 60.7 1.5 kJ/mol.
| Method | ||||||||
| CCSD(T) | post-CCSD(T) | |||||||
| Slopes | Basis Set | MP2 | CCSD | vvTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -6.53 | -4.74 | -5.58 | -5.47 | -5.35 | -5.36 | -5.49 |
| 13 | DZ | -6.99 | -4.89 | -5.81 | -5.69 | -5.56 | -5.54 | -5.70 |
| 14 | DZ | -7.44 | -5.08 | -6.07 | -5.95 | -5.81 | -5.81 | -5.97 |
| 17 | TZ | -9.93 | -6.38 | -7.96 | -7.77 | -7.57∗ | -7.57∗ | -7.79∗ |
| 12 | {T,Q}Z | -9.53 | -6.55 | -8.11 | -7.97 | -7.81∗ | -7.82∗ | -8.00∗ |
| 14 | {T,Q}Z | -10.58 | -6.96 | -8.73 | -8.56 | -8.36∗ | -8.37∗ | -8.59∗ |
| 17 | {T,Q}Z | -9.03 | -8.84 | -8.64∗ | -8.64∗ | -8.86∗ | ||
| 12 | a{Q,5}Z | -9.54 | -6.55 | -8.04 | -7.98 | -7.82∗ | -7.82∗ | -8.01∗ |
| 14 | a{Q,5}Z | -8.63 | -8.47 | -8.28∗ | -8.26∗ | -8.50∗ | ||
| 17 | a{Q,5}Z | -8.94 | -8.74 | -8.55∗ | -8.55∗ | -8.77∗ | ||
@@
| Method | ||||||||
| CCSD(T) | post-CCSD(T) | |||||||
| Slopes | Basis Set | MP2 | CCSD | vvTight | canonical | CCSDT-2 | CCSDT | CCSDT(Q) |
| 12 | DZ | -7.03 | -4.87 | -5.93 | -5.72 | -5.59 | -5.59 | -5.72 |
| 13 | DZ | -7.60 | -5.15 | -6.24 | -6.09 | -5.94 | -5.92 | -6.09 |
| 14 | DZ | -8.14 | -5.44 | -6.62 | -6.46 | -6.30 | -6.29 | -6.46 |
| 17 | TZ | -10.71 | -6.94 | -8.62 | -8.46 | -8.24∗ | -8.23∗ | -8.46∗ |
| 12 | {T,Q}Z | -9.65 | -6.57 | -8.15 | -8.01 | -7.83∗ | -7.82∗ | -8.01∗ |
| 14 | {T,Q}Z | -10.68 | -6.98 | -8.75 | -8.59 | -8.37∗ | -8.36∗ | -8.59∗ |
| 17 | {T,Q}Z | -9.01 | -8.85 | -8.63∗ | -8.62∗ | -8.86∗ | ||
| 12 | a{Q,5}Z | -9.58 | -6.53 | -8.04 | -7.95 | -7.77∗ | -7.77∗ | -7.96∗ |
| 14 | a{Q,5}Z | -8.59 | -8.43 | -8.20∗ | -8.20∗ | -8.43∗ | ||
| 17 | a{Q,5}Z | -8.85 | -8.69 | -8.47∗ | -8.46∗ | -8.69∗ | ||
@@
An analogous estimation is performed for the ethylene stack series in Table 4 and 5. To extrapolate CCSDT and CCSDT(Q) values, we utilize the difference between CCSDT and CCSD(T) or between CCSDT(Q) and CCSD(T) results obtained with the DZ basis set. The most accurate CCSD(T) calculations were performed using the TZ basis set for the entire ethylene stack series (ethylene, tetraene, pentaene up to tetradecaneheptaene) and the QZ basis set for ethylene–decanepentaene. These results were subsequently used to extrapolate values to larger basis sets. Interestingly, canonical CCSD(T) even slightly underestimates the correlation interaction energy slopes relative to CCSDT(Q), showing that the correlation slopes obtained are rather dependent on the systems chosen. Overall, the post-CCSD(T) contributions are also smaller for the ethylene stacks, as the CCSDTCCSD(T) difference amounts to 2.7% of the correlation contribution to the interaction energy. For the acene series, the CCSDTCCSD(T) difference is more substantial at 4.6% — presumably because of the aromatic rings.
The more economical CCSD(T)Λ method yields qualitatively the same answer as CCSDT-2, but not quantitatively.